Achondroplasia is a dominant condition that leads to dwarfism. Hereditary deafness is a recessive trait. John,...

  1. Achondroplasia is a dominant condition that leads to dwarfism. Hereditary deafness is a recessive trait. John, who is achondroplastic and hears normally, marries Ann, who has normal height and hearing, but whose father was deaf. They have a non-achondroplastic 3 year old son who is deaf.
    1. Draw the pedigree for this family, indicating all the known family members.
    2. Give the genotypes of John, Ann, and their son.
    3. If John and Ann have another child, what is the probability that
      1. The child is achondroplastic?
      2. The child is deaf?
      3. The child is both achondroplastic and deaf?
  1. In German cockroaches, curved wing (cv) is recessive to normal wing (cv+), and black body (b) is recessive to brown body (b+). The wing shape and body color genes assort independently. A cockroach has the genotype cv+cv b+b. Draw arrangements of chromosomes during metaphase I of meiosis to show independent assortment of the above alleles, clearly indicating the gene loci on the chromosomes.
Answer #1

aaHh × AaHh

Gametes aH ah
AH AaHH (achondroplasia, normal hearing) AaHh (achondroplasia, normal hearing)
Ah AaHh (achondroplasia, normal hearing) Aahh (achondroplasia, deaf)
aH aaHH (normal height, normal hearing) aaHh (normal height, normal hearing)
ah aaHh (normal height, normal hearing) aahh (normal height, deaf)

Total number of children = 8

Number of achondroplastic children = 4

Probability of achondroplastic children = 4/8 = 1/2

Number of deaf children = 2

Probability of deaf children = 2/8 = 1/4

Number of children with both achondroplasia and deafness = 1

Probability of children with both achondroplasia and deafness = 1/8

Please rate.